OpenStudy (phebe):

Solve the equation by completeing the square. Round to the nearest 10th. A- 1.1,9.1 B- 1.1,-9.1 C- -1.1,9.1 D- -1.1,-9.1

OpenStudy (phebe):

@mjoxprm360

OpenStudy (anonymous):

whats the equation?

OpenStudy (phebe):

hold on k

OpenStudy (anonymous):

ok :)

OpenStudy (phebe):

i mean x^2+8x=10

OpenStudy (phebe):

anybody gonna help me wit dizz question

OpenStudy (anonymous):

im working on it

OpenStudy (phebe):

ok kcool sorry buggin u

OpenStudy (anonymous):

is this the equation? x² + 8x + 10 = 0

OpenStudy (anonymous):

your not bugging me :)

OpenStudy (phebe):

no its x^2+8x=10

OpenStudy (phebe):

ohh ok kcool @mjoxprm360

OpenStudy (anonymous):

ok

OpenStudy (phebe):

u still working on it @mjoxprm360

OpenStudy (phebe):

@skullpatrol

OpenStudy (phebe):

aswerr ittttt XD so i can finish my work @mjoxprm360

OpenStudy (anonymous):

ok

OpenStudy (phebe):

thanx

OpenStudy (phebe):

@Luigi0210

OpenStudy (luigi0210):

So you have \( x^2+8x=10\), you have to get it into standard form, so subtract 10 from both sides and you'll get \( x^2+8x-10=0\). Now just factor.

OpenStudy (luigi0210):

Use the quadratic formula actually.

OpenStudy (phebe):

ok than so wat u think is the answer @Luigi0210

OpenStudy (luigi0210):

I know the answer, but I can't give it out directly >.<

OpenStudy (phebe):

well foreal but im in a hurry